OpenStudy (mendicant_bias):

Going over some basic Differential Equations. Posted below momentarily.

OpenStudy (mendicant_bias):

Solve the differential equation.\[\frac{y}{\sqrt{4-x^2}}=y^2+5\]

OpenStudy (mendicant_bias):

@dan815 , could you help me out with this? I really don't remember how to do this...

OpenStudy (anonymous):

Is that \(y'\) on the left?

OpenStudy (mendicant_bias):

Sorry, I forgot an entire term; the lefthand side is multiplied by y prime.

OpenStudy (anonymous):

No problem. The equation is separable: \[\frac{y}{\sqrt{4-x^2}}y'=y^2+5\] Separating yields \[\frac{y}{y^2+5}dy=\sqrt{4-x^2}~dx\]

OpenStudy (anonymous):

The left can be integrated with a substitution, the right with a trigo sub.

OpenStudy (mendicant_bias):

Alright, gonna take a shot at the rest from here, one sec. Thank you so far.

OpenStudy (anonymous):

yw

OpenStudy (mendicant_bias):

\[u = y^2+5, \ du = 2y \ dy; \ \ \ \]

OpenStudy (mendicant_bias):

Ugh, trig subs....is this the thing where.......one sec, lol. Gotta think about trig identities for a second again.

OpenStudy (mendicant_bias):

Could x be substituted for either 2sin(theta) or 2cos(theta), if I recall correctly?

OpenStudy (anonymous):

In this case, yes, either works. The sine sub is usually taken to avoid working with the pesky negative the cosine gives.

OpenStudy (mendicant_bias):

Hey, I'm not too familiar with LaTeX, do you know how I "turn off" the \rm statement in the middle of a sentence? e.g. I want to type upright in text using it, but then I want to put a math formula in the sentence. I'll do it either way, but I want to figure that out.

OpenStudy (freckles):

\text{ }

OpenStudy (freckles):

i don't know maybe i misunderstood

OpenStudy (mendicant_bias):

\[\frac{y}{y^2 +5}\frac{dy}{dx}=\sqrt{4-x^2}; \ \ \ \frac{1}{2}\int\limits_{}^{}\frac{du}{u}du=\int\limits_{}^{}\sqrt{4-4 \sin(\theta)}dx\]

OpenStudy (anonymous):

I think OP is wondering how to type something like \(\sum\) in-line with text? In which case you'd wrap your code in backslash-parentheses, as in `\( <code> \)`.

OpenStudy (mendicant_bias):

Let me try it out. \[\rm Let \( swag \)\] (Lol, didn't work)

OpenStudy (freckles):

something like this: \[\frac{ \text{ see a fish }}{ \text{ be a fish } } \text{ is a fraction }\]

OpenStudy (mendicant_bias):

But yeah, I want math formulas with slanted, italicized variables and all, in the middle of a sentence in LaTeX.

OpenStudy (mendicant_bias):

\[\frac{1}{2}\ln(y^2+5)+c_{1}=\int\limits_{}^{}2\cos(\theta)dx\] I forget from here how you change the right-hand's variable of integration to d-theta in a situation where you use trig substitution.

OpenStudy (mendicant_bias):

AHH yeah, I got it, it was because I was using \rm, I know you mentioned \text{}, but I've never used it, thank you~!

OpenStudy (anonymous):

Maybe like this?\[\text{Let }swag\] A few ways you can do it. `\[\text{Let }swag\]` (as above) `\[\text{Let }\emph{swag}\]` (as below) \[\text{Let }\textit{swag}\] The differences are slight, mainly with the kerning and scaling of the letters.

OpenStudy (anonymous):

Sorry, `\emph` doesn't work on this site... I actually had used `\textit{...}`.

OpenStudy (mendicant_bias):

(Lol) Alright, but yeah, moving forward with the problem, how do I deal with that differential being dx? I forgot how this works.

OpenStudy (freckles):

http://openstudy.com/study#/groups/LaTeX%20Practicing!%20%3A) this is a good group to learn latex in people have made a lot of threads on latex tutorials

OpenStudy (anonymous):

You set \(x=2\sin\theta\), which gives what differential for \(x\)? \(dx=\cdots\)

OpenStudy (mendicant_bias):

Oh, derp. Yeah, I gotchu.

OpenStudy (mendicant_bias):

>The fact that I didn't realize that feelsbadman.jpg \[\frac{1}{2}\ln(y^2+5)+c_{1}=\int\limits_{}^{}2\cos(\theta)[2\cos(\theta)d \theta]\]

OpenStudy (mendicant_bias):

I'm moving slowly on this because the whole form of this thing looks like it's about to play a trick on me, the integrand being identical to the differential element, lol....alright. Integrating the RHS:

OpenStudy (mendicant_bias):

\[4\int\limits_{}^{}\cos^{2}(\theta)d \theta = 2\int\limits_{}^{}(1+\cos(2 \theta)d \theta\]

OpenStudy (mendicant_bias):

\[2 \int\limits_{}^{}1+\cos(2 \theta))d \theta = 2\bigg[\theta+\frac{\sin(2 \theta)}{2}\bigg]\]

OpenStudy (mendicant_bias):

\[2 \theta + \sin(2 \theta)+c_{2}=\frac{1}{2}\ln(y^2+5)+c_{1}\]

OpenStudy (mendicant_bias):

(Sloppy introduction of c_{2}, whoops.)

OpenStudy (mendicant_bias):

Uhh, something subbing x stuff where theta is.

OpenStudy (freckles):

I like to draw a right triangle for my trib substitution and use it as a reference later to put back in terms of what I started with

OpenStudy (freckles):

also sin(2u)=2sin(u)cos(u)

OpenStudy (mendicant_bias):

Yeah, I'm just being a dummy and forgetting how to do it algebraically as well.

OpenStudy (mendicant_bias):

One sec.

OpenStudy (mendicant_bias):

\[2\sin(\theta) = x; \ \ \ \sin(\theta) = \frac{x}{2}.\] Can set up that reference triangle from there, one moment.

OpenStudy (mendicant_bias):



OpenStudy (freckles):

you only really need that triangle for the cos(theta) part

OpenStudy (mendicant_bias):

:| I didn't need it for the cos(theta) part ironically (personally)-but I can see it being useful if I got a spaghetti mess of different trig functions from the original. Alright, either way, need to use the identity you mentioned.

OpenStudy (mendicant_bias):

\[2\theta + \sin(2 \theta) = 2\sin (\theta) \cos (\theta)\]

OpenStudy (freckles):

\[2 \theta+\sin(2 \theta)=2 \theta+2 \sin(\theta)\cos(\theta)\]

OpenStudy (mendicant_bias):

Theta by itself is:\[\sin^{-1}(x/2)\]

OpenStudy (mendicant_bias):

This is looking stupidly more complex than I expected for just a run-of-the-mill DE, I sure hope I didn't make a mistake somewhere.

OpenStudy (mendicant_bias):

\[\sin^{-1}(x/2)+x\frac{\sqrt{4-x^2}}{2}\]

OpenStudy (mendicant_bias):

\[\sin^{-1}(x/2)+\frac{x \sqrt{4-x^2}}{2}=\frac{1}{2}\ln(y^{2}+5)+C\] (Just gonna collect together those arbitrary constants into one, I think I'm allowed to do that, but not 100% sure.)

OpenStudy (mendicant_bias):

Now what do I do?

OpenStudy (freckles):

that looks good

OpenStudy (mendicant_bias):

Well, don't I have to solve for something, or....something? Lol. I don't remember how to solve DE's like this, but I feel like I have to express the solution in some format other than this.

OpenStudy (mendicant_bias):

@wio , do I stop here, or does this look good?

OpenStudy (freckles):

I think that form is find to express the answer but you can try to solve for y just so you know it isn't always easy or possible to solve for y

OpenStudy (freckles):

fine*

OpenStudy (mendicant_bias):

Yeah, that's why, I feel quite a bit like (this was a sample final exam question) we would have to very explicitly express y as a function of x for the solution.